Grade 12th passPhysical Chemistry

The radius of La3+ (Atomic number of La=57) is 1.06A. Which one of the following given values will be closest to the radius of Lu3+ (Atomic number of Lu = 71)?
(1) 40 A
(2) 1.06 A
(3) 0.85 A
(4) 1.60 A

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer ::

Radius, r ∝ 1/Z
r1/r2 = Z2/Z1
1.06/r2 = 71/57
So r2 = 1.06×57/71 = 0.85
Hence option (3) is the answer.


Thanks